Experiment 12 THE DETERMINATION OF THE NEUTRALIZING ABILITY OF ANTACIDS The Pre-Lab 1. A 1.032 gram sample of an antacid was this is my prelab for class, I'm really struggling to understand so if someone could help me break down the problems that would be nice
Answer #1

Information provided is listed as:

Antacid taken = 1.032g in 50 mL water

HCl added = 49.33 mL of 0.488 M concentration

After neutralizing the antacid, NaOH required to neutralize the remained HCl = 29.33 mL of 0.510 M NaOH

a). The source of the two bases that react with HCl is antacid and Sodium hydroxide.

b). No. of moles of acid used in experiment = concentration (M) * volume (in L)

= 0.488 M * (49.33/1000) = 0.024 moles

c). total no. of moles of HCl that must be neutralized by antacid and NaOH is 0.024 moles

d). No. of moles of NaOH added = concentration (M) * volume (in L)

= 0.510* (29.33/1000) = 0.015 moles